Due ne wed 30A soccer hall is kicked from the ground with an initial speed of 19.5 mvs at an upward angle of 45.A player 55 m away in the direction of the kick starts running to moet the ball at that in slant. What imust be his average speed if he is to moet the ball just helore it hits the ground? ch4 15 SSM ILW A particle leaves the origin with an initial vchK- ity . (3.00i) mrs ind a constant acceleration 급 ( 1.00i 0.500)) m/s. When it reaches its maximum z courdinate, whal are its (a) velocity and (b) position veetor? 47 SSM www A batter hits a pitched ball when the center of the ball is 1.22 m above the ground. The ball leaves the bat al an +25 ; Thc current world-record motorcycle jump ts 77.0 m angle of 45 th the ground with that launch-the hal should have set by Jason Renic. Assume that he left the take-off ramp at a horizontal range (returning to the launch level) of 107 m. (a) 12.0P to the horizontal and that the take-off and landing Does the ball clear a 7.32-m-high fence that is 97.5 m horizontally hcights are the same. Neglocting air drag. determine his take-off from the launch point? (b) At the fence, what is the distance he- speed twcen the fence top and the ball center?
Answer #1

Time of flight = 2 u sin theta /g = 2*19.5 sin 45 degree /9.8

= 2.814s

Range = u^2 sin 2 theta/g = 19.5^2 sin 90 degree /9.8

= 38.8 m

Average speed of his run = (55-38.8)/2.814

= 5.76 m/s answer